9 is what percent of 660
Answer by redjenna333(3)

Use the equation
A/b=p/100
A is the part you are solving for
B is the whole number
P is the percent of 100
so
A/660=9/100
cross multiply for 100a=5,940
then divide by 100 for a
so a=59.4
and you can check by dividing 59.4 by 660 it comes out as 9%